OpenStudy (anonymous):

1 3w 4 - + - = - 5 15 5 Show how you get it and a check please

OpenStudy (anonymous):

3w/15 = w/5 so 1/5+w/5 = (1+w)/5 (1+w)/5 = 4/5 so w = 3

OpenStudy (anonymous):

check: 1/5 +9/15 = 4/5 1/5 +3/5 = 4/5 4/5=4/5